OpenStudy (anonymous):

Find a quadratic equation in standard form with integral coefficients and having numbers as roots.

OpenStudy (anonymous):

\[\frac{ -5-\sqrt{2} }{ 3 } , \frac{ -5+\sqrt{2} }{ 3 }\]

OpenStudy (tkhunny):

Well, do it. If the problem were 3, 4, what would you do?

OpenStudy (anonymous):

I would find b and c, but I can't simplify the i in the end

OpenStudy (tkhunny):

3,4 is like this y = (x-3)(x-4) = x^2 - 7x + 12 2/3, 3/4 is like this y = (x - 2/3)(x - 3/4) = x^2 - (17/12)x + (1/2) Or maybe y = (3x-2)(4x-3) = 12x^2 -17x + 6 \(2+\sqrt{5}\), \(2-\sqrt{5}\) is like this \(y = (x - (2+\sqrt{5}))(x - (2-\sqrt{5}))\) = ??? That's enough of that. You finish taht one and then let's ahve a look at your original problem statement.